-2 (-5 w - x) + (2 x + 2 w)
Answer by lenny460(1073)

-2(-5w - x) + (2x + 2w)

-2 * -5w = 10w
-2 * -x = 2x

(10w + 2x) + (2x + 2w)
Remove the parentheses:
10w + 2x + 2x + 2w
Combine like terms:
10w + 2w = 12w
2x + 2x = 4x

Answer:
12w + 4x









Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com